# Rambo: RL-augmented Model-based Whole-body Control for Loco-manipulation

### v2

@@ V Results
Forward velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{x}$ $[-0.5,0.5]$ (m/s)
Side velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{y}$ $[-0.5,0.5]$ / $[-0.0,0.0]$ (m/s)
turning velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\omega}_{z}$ $[-0.5,0.5]$ (rad/s)
EE position $x$ $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}^{m}_{i}})_{x}$ $[0.1934,0.5]$ / $[0.15,0.30]$ (m)
EE position $y$ $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}^{m}})_{y}$ $[0,0.2]$ (m)
EE position $z$ $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}^{m}})_{z}$ $[0.0,0.4]$ / $[0.3,0.9]$ (m)
EE force in $x,y,z$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{u}}$ $[-30,30]$ / $[-20,20]$ (N)
@@
Add mass for base $[-2.0,2.0]$ (kg)
CoM offset for base $[-0.05,0.05]$ (m)
@@
Base position noise $\bm{p}_{x},\bm{p}_{y}$ $[-0.05,0.05]$ (m)
Base position noise $\bm{p}_{z}$ $[-0.02,0.02]$ (m)
Base orientation noise $\bm{R}$ $[-0.05,0.05]$
Base linear velocity noise $\bm{v}$ $[-0.1,0.1]$ (m/s)
Base angular velocity noise $\bm{\omega}$ $[-0.15,0.15]$ (m/s)
Joint position noise $\bm{q}_{j}$ $[-0.01,0.01]$ (rad)
Joint velocity noise $\dot{\bm{q}}_{j}$ $[-1.5,1.5]$ (rad/s)
TABLE IV: Detailed command sampling and domain randomization. ()/()($\cdot$)/($\cdot$)( ⋅ ) / ( ⋅ ) represents different values used for quadruped and biped tasks respectively. The orientation noise is estimated from the noise sampled on unit quaternion representations. Besides the terms listed above, we also add random push on the robot base within an episode. Note that the EE position and forces are listed only for the front-left limb of the robot. For biped tasks, we sample the quantities in symmetry for the front-right limb.
@@ V-A Quantitative Evaluation
  • Rambo-base: WBC with a feedback policy outputting acceleration correction $\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{a}}}$ only;
@@
Rambo-joint $0.108\pm 0.014$ $0.101\pm 0.032$ $0.046\pm 0.007$ $0.306\pm 0.046$ $0.383\pm 0.109$ $0.134\pm 0.044$
Rambo-ff $0.374\pm 0.036$ $0.404\pm 0.154$ $0.286\pm 0.048$ $0.320\pm 0.061$ $0.389\pm 0.187$ $0.447\pm 0.026$
Rambo $\mathbf{0.087\pm 0.009}$ $\mathbf{0.085\pm 0.022}$ $\mathbf{0.039\pm 0.003}$ $0.286\pm 0.039$ $0.352\pm 0.075$ $\mathbf{0.036\pm 0.002}$
@@ V-C Further Comparison with Vanilla Policy
We attribute this discrepancy to the relatively large exploration space in the vanilla training setup, which makes it difficult for the policy to consistently prioritize accurate EE tracking while generating appropriate contact forces through its end-effectors. Instead, Rambo separates feedforward torque from WBC and joint feedback from policy, which allows the learned policy to provide corrective offsets relative to reference positions with a much smaller exploration space, improving robustness without significantly altering the desired kinematic motion.
@@ VI Conclusion
Rambo is not without limitations. Currently, it relies solely on proprioceptive information, which negatively affects performance due to drift in state estimation. As a next step, we aim to incorporate additional sensing modalities to enhance robustness and accuracy. Nevertheless, we see strong potential for Rambo in future loco-manipulation research, including the integration of full-order WBC dynamics models and its application to higher degree-of-freedom humanoids. Other promising directions include extending the framework with online model adaptation to further improve generalization and precision.

### v4

@@ V Results
Forward velocity ${{}_{\mathcal{P}}}\hat{v}_{x}$ [0.5,0.5][-0.5,0.5] (m/s)
Side velocity ${{}_{\mathcal{P}}}\hat{v}_{y}$ [0.5,0.5][-0.5,0.5] / [0.0,0.0][-0.0,0.0] (m/s)
Turning velocity ${{}_{\mathcal{P}}}\hat{\omega}_{z}$ [0.5,0.5][-0.5,0.5] (rad/s)
EE position $x$ $({{}_{\mathcal{P}}}{\hat{\bm{p}}^{m}_{i}})_{x}$ $[0.1934,0.5]$ / $[0.15,0.30]$ (m)
EE position $y$ $({{}_{\mathcal{P}}}{\hat{\bm{p}}^{m}})_{y}$ $[0,0.2]$ (m)
EE position $z$ $({{}_{\mathcal{P}}}{\hat{\bm{p}}^{m}})_{z}$ $[0.0,0.4]$ / $[0.3,0.9]$ (m)
EE force in $x,y,z$ ${{}_{\mathcal{P}}}\hat{\bm{u}}$ [30,30][-30,30] / [20,20][-20,20] (N)
@@
Add mass for base [2.0,2.0][-2.0,2.0] (kg)
CoM offset for base [0.05,0.05][-0.05,0.05] (m)
@@
Base position noise $\bm{p}_{x},\bm{p}_{y}$ [0.05,0.05][-0.05,0.05] (m)
Base position noise $\bm{p}_{z}$ [0.02,0.02][-0.02,0.02] (m)
Base orientation noise $\bm{R}$ [0.05,0.05][-0.05,0.05]
Base linear velocity noise $\bm{v}$ [0.1,0.1][-0.1,0.1] (m/s)
Base angular velocity noise $\bm{\omega}$ [0.15,0.15][-0.15,0.15] (m/s)
Joint position noise $\bm{q}_{j}$ [0.01,0.01][-0.01,0.01] (rad)
Joint velocity noise $\dot{\bm{q}}_{j}$ [1.5,1.5][-1.5,1.5] (rad/s)
TABLE IV: Detailed command sampling and domain randomization. $(\cdot)/(\cdot)$ represents different values used for quadruped and biped tasks respectively. The orientation noise is estimated from the noise sampled on unit quaternion representations. Besides the terms listed above, we also add random push on the robot base within an episode. Note that the EE position and forces are listed only for the front-left limb of the robot. For biped tasks, we sample the quantities in symmetry for the front-right limb.
@@ V-A Quantitative Evaluation
  • Rambo-base: WBC with a feedback policy outputting acceleration correction $\Delta{{}_{\mathcal{P}}}{\hat{\bm{a}}}$ only;
@@
Rambo-joint $8\pm 0.014 0.101±0.0$$1\pm 0.032$ $0.046\pm 0.007$ $0.306\pm 0.046$ $0.383\pm 0.109$ $0.134\pm 0.044$
Rambo-$f$ $0.374\pm 0.036$ $0.404\pm 0.154$ $0.286\pm 0.048$ $0.320\pm 0.061$ $0.389\pm 0.187$ $0.447\pm 0.026$
[…]
@@ V-C Further Comparison with Vanilla Policy
We attribute this discrepancy to the relatively large exploration space in the vanilla training setup, which makes it difficult for the policy to consistently prioritize accurate EE tracking while generating appropriate contact forces through its end-effectors. Instead, Rambo separates feedforward torque from WBC and joint feedback from policy, which allows the learned policy to provide corrective offsets relative to reference positions with a much smaller exploration space, improving robustness without significantly altering the desired motion.
@@ VI Conclusion
Rambo is not without limitations. Currently, it relies solely on proprioceptive information, which negatively affects performance due to drift in state estimation. As a next step, we aim to incorporate additional sensing modalities to enhance robustness and accuracy. Nevertheless, we see strong potential for Rambo in future loco-manipulation research, including the integration of full-order WBC dynamics models and its application to humanoids. Other promising directions include extending the framework with online model adaptation to further improve generalization and precision.
VII Acknowledgment
We thank Zijun Hui, Taerim Yoon for developing the hardware testing software stack. We thank Yu (Antares) Zhang for helping the hardware experiments.